The sum of the first 3 terms is 42 and the sum of the next 3 terms is 5 1/4. Determine the first 3 terms of geometric progression.
Answer by richwmiller(17219)

S=t*(1 - r^n)/(1 - r)
42=t*(1 - r^3)/(1 - r),
47.25=t*(1 - r^6)/(1 - r)
r=0.5, t=24
24.12,6 first 3 terms
3, 1.5, .75 second three terms